# A2G-QFL: Adaptive Aggregation with Two Gains in Quantum Federated learning

### v1

@@ I Introduction
Federated learning (FL) [mcmahan2017communication, nanayakkara2024understanding] deployed over quantum-enabled and heterogeneous classical networks exhibits several fundamental challenges. These include non-uniform client reliability, stochastic variations in teleportation fidelity, device-level instability, circular or manifold-valued model parameters (e.g., rotation angles in variational quantum circuits), and geometry-induced drift among local model trajectories.
@@ I-A Limitations of Prior Work
FedAvg does not account for geometric considerations or network QoS. Riemannian FL lacks integration of latency and fidelity metrics. Quantum FL research primarily addresses quantum model development rather than aggregation strategies [liu2025practical]. Federated learning has traditionally relied on euclidean parameter averaging, as exemplified by FedAvg. However, FedAvg assumes euclidean averaging and demonstrates poor performance in non-independent and identically distributed (non-IID) settings. This approach, as described by McMahan et al. [mcmahan2017communication], implicitly assumes that local model updates occur within a locally linear region of the loss landscape. Extensive empirical and theoretical analyses indicate that statistical heterogeneity, device variability, and client drift result in highly non-linear and curved parameter trajectories. These factors render Euclidean averaging unstable and frequently divergent in heterogeneous environments [karimireddy2020scaffold, li2020federated, wang2020tackling]. This development has led to an increasing body of research on Riemannian and geometry-aware federated optimization, in which model parameters are represented as points on smooth manifolds and aggregation is conducted using exponential and logarithmic maps [bonnabel2013stochastic, sakai2021riemannian, zhang2016riemannian, hsieh2023riemannian, zhang2024nonconvex].
The geometric challenges are even more pronounced in quantum federated learning (QFL), where variational quantum circuits parameterize unitary transformations via angles living on periodic manifolds [mitarai2018quantum, schuld2019quantum, cerezo2021variational, abbas2021power, khan2025dt]. These parameters evolve under the influence of hardware noise, decoherence, shot-induced gradient variance, and in distributed implementations teleportation fidelity fluctuations, all of which distort the effective optimization landscape and make naive euclidean aggregation misaligned with the underlying parameter geometry [beer2020training, huang2025vast]. Prior QFL studies either disregard geometric structure or address it heuristically, resulting in unstable global updates and high variance under quantum channel imperfections [chen2021federated, qi2023optimizing]. However, A2G provides a general dual-gain aggregation framework for mixed classical–quantum FL, whereas FAAA-$\beta$ which we proposed in our previous work is a specialized teleportation-conditioned angular consensus rule for variational quantum models operating on toroidal parameter spaces.
@@ VI Discussion
In this work we deliberately adopt a geometry–gain update rather than a gradient-based server update. This choice is principled for both classical and quantum FL. In variational quantum circuits, server-side gradients are notoriously expensive and fragile due to sampling noise and barren-plateau effects [cerezo2021variational], and parameter-shift evaluations incur a substantial communication and runtime overhead when aggregated across many clients [schuld2019quantum]. These issues are exacerbated when model updates must be transmitted over noisy or teleportation-based links in QFL [li2020federated]. By contrast, client-side optimisers such as SPSA can estimate local gradients with low communication cost and are already widely used in noisy quantum settings. Restricting the server to a geometry-controlled interpolation step therefore avoids expensive global gradient computation while still steering the model towards consensus.
The geometry gain parameter $\beta$ provides a stable, communication-efficient, gradient-free mechanism to blend the previous global model with the QoS-weighted aggregate. Small but nonzero values (e.g., $\beta=0.05$ on Breast-Lesions-USG) consistently outperform the FedAvg-style setting ($\beta=1.0$) by more than ten percentage points in best and late-round accuracy, indicating that geometry-aware damping can mitigate both client drift and QoS-induced bias. This behaviour is consistent with observations from classical FL, where momentum, EMA-style smoothing, and proximal or control-variate corrections are used to stabilise aggregation under heterogeneity [mcmahan2017communication, karimireddy2020scaffold, li2020federated]. Our results therefore suggest that geometry-gain A2G offers a practically attractive and theoretically grounded alternative to gradient-based server updates in both quantum and hybrid quantum–classical federated deployments.

### v2

@@ I Introduction
Federated learning (FL) [12, 14] deployed over quantum-enabled and heterogeneous classical networks exhibits several fundamental challenges. These include non-uniform client reliability, stochastic variations in teleportation fidelity, device-level instability, circular or manifold-valued model parameters (e.g., rotation angles in variational quantum circuits), and geometry-induced drift among local model trajectories.
@@ I-A Limitations of Prior Work
FedAvg does not account for geometric considerations or network QoS. Riemannian FL lacks integration of latency and fidelity metrics. Quantum FL research primarily addresses quantum model development rather than aggregation strategies [11]. Federated learning has traditionally relied on euclidean parameter averaging, as exemplified by FedAvg. However, FedAvg assumes euclidean averaging and demonstrates poor performance in non-independent and identically distributed (non-IID) settings. This approach, as described by McMahan et al. [12], implicitly assumes that local model updates occur within a locally linear region of the loss landscape. Extensive empirical and theoretical analyses indicate that statistical heterogeneity, device variability, and client drift result in highly non-linear and curved parameter trajectories. These factors render Euclidean averaging unstable and frequently divergent in heterogeneous environments [8, 10, 18]. This development has led to an increasing body of research on Riemannian and geometry-aware federated optimization, in which model parameters are represented as points on smooth manifolds and aggregation is conducted using exponential and logarithmic maps [3, 16, 19, 6, 20].
The geometric challenges are even more pronounced in quantum federated learning (QFL), where variational quantum circuits parameterize unitary transformations via angles living on periodic manifolds [13, 17, 4, 1, 9]. These parameters evolve under the influence of hardware noise, decoherence, shot-induced gradient variance, and in distributed implementations teleportation fidelity fluctuations, all of which distort the effective optimization landscape and make naive euclidean aggregation misaligned with the underlying parameter geometry [2, 7]. Prior QFL studies either disregard geometric structure or address it heuristically, resulting in unstable global updates and high variance under quantum channel imperfections [5, 15]. However, A2G provides a general dual-gain aggregation framework for mixed classical–quantum FL, whereas FAAA-$\beta$ which we proposed in our previous work is a specialized teleportation-conditioned angular consensus rule for variational quantum models operating on toroidal parameter spaces.
@@ VI Discussion
In this work we deliberately adopt a geometry–gain update rather than a gradient-based server update. This choice is principled for both classical and quantum FL. In variational quantum circuits, server-side gradients are notoriously expensive and fragile due to sampling noise and barren-plateau effects [4], and parameter-shift evaluations incur a substantial communication and runtime overhead when aggregated across many clients [17]. These issues are exacerbated when model updates must be transmitted over noisy or teleportation-based links in QFL [10]. By contrast, client-side optimisers such as SPSA can estimate local gradients with low communication cost and are already widely used in noisy quantum settings. Restricting the server to a geometry-controlled interpolation step therefore avoids expensive global gradient computation while still steering the model towards consensus.
The geometry gain parameter $\beta$ provides a stable, communication-efficient, gradient-free mechanism to blend the previous global model with the QoS-weighted aggregate. Small but nonzero values (e.g., $\beta=0.05$ on Breast-Lesions-USG) consistently outperform the FedAvg-style setting ($\beta=1.0$) by more than ten percentage points in best and late-round accuracy, indicating that geometry-aware damping can mitigate both client drift and QoS-induced bias. This behaviour is consistent with observations from classical FL, where momentum, EMA-style smoothing, and proximal or control-variate corrections are used to stabilise aggregation under heterogeneity [12, 8, 10]. Our results therefore suggest that geometry-gain A2G offers a practically attractive and theoretically grounded alternative to gradient-based server updates in both quantum and hybrid quantum–classical federated deployments.